OpenStudy (anonymous):

Find the common ratio of geometric sequence with a1=3/5 and a6=1875/1024

OpenStudy (anonymous):

Using the nth-term formula for a geometric series,\[a_6=a_1r^5\]We know that \(a_6 = \frac{1875}{1024}\) and \(a_1=\frac 35\) from the problem statement. So, now we just plug it in and solve for \(r\). Can you solve it from there?